### 01 Apr 2026 v2

# Cognitive Friction: A Decision-Theoretic Framework for Bounded Deliberation in Tool-Using Agents

Abstract
Autonomous tool-using agents operating in networked environments must decide both which information source to query and when to stop querying and act. Without principled bounds on information-acquisition costs, unconstrained agents exhibit systematic failure modes: excessive tool use under congestion, prolonged deliberation under time decay, and brittle behavior under ambiguous evidence. We propose the Triadic Cognitive Architecture (TCA), a unified decision-theoretic framework that formalizes these failure modes through the concept of Cognitive Friction. By synthesizing nonlinear filtering theory, congestion-dependent cost dynamics, and HJB optimal stopping, we model deliberation as a stochastic control problem over a joint belief–congestion state space, where information acquisition is explicitly priced by tool-dependent signal quality and live network load. Rather than relying on arbitrary heuristic stop-tokens or fixed query budgets, TCA derives an HJB-inspired stopping boundary and instantiates a computable rollout-based approximation of belief-dependent value-of-information with a net-utility halting condition. We validate the framework on two controlled simulation environments, the Emergency Medical Diagnostic Grid (EMDG) and the Network Security Triage Grid (NSTG), designed to isolate the decision-theoretic quantities of interest (stopping quality, action selection under congestion, and temporal urgency) under reproducible conditions. TCA reduces time-to-action while improving resource outcomes without degrading accuracy: over greedy baselines, TCA gains 36 viability points in EMDG and 33 integrity points in NSTG. Two purposive stopping baselines confirm that the joint optimization of selection and stopping is essential; stopping rules alone recover at most 4 viability points. A univariate sensitivity sweep over $\alpha$, $\beta$, and $\lambda_{S}$ shows stable accuracy and interpretable tradeoffs, and an empirical sweep over $\eta\in\{0,0.1,0.3,0.5\}$ confirms $\eta{=}0$ is optimal on EMDG trajectories under high temporal urgency.
1 Introduction
Autonomous agents that interact with networked environments through tool queries (whether in medical triage, network security, or open-ended task execution) face a fundamental decision problem: which information source to query next, and when to stop querying and act. Iterative query-then-decide loops such as ReAct yao2023react , Tree-of-Thoughts yao2023tot , and AutoGPT autogpt2023 have demonstrated that sequential querying substantially improves decision quality. However, these systems operate under the implicit assumption of a frictionless cognitive environment: gathering information is presumed to be instantaneous, structurally free, and uniformly reliable.
When deployed outside of static text-generation benchmarks and into dynamic, physical, or highly networked environments, this assumption mathematically can lead to systematic failure modes. Without intrinsic bounds on spatial topology, temporal pacing, or epistemic resolution, unconstrained agents exhibit three critical failure modes:
  1. 1.
    Congestion Saturation (Space): Indiscriminate API querying cascades into network congestion and rising per-query load costs, as agents fail to price the structural cost of routing information through a multi-agent swarm or distributed database.
  2. 2.
    Infinite Deliberation (Time): Lacking a continuous-time pacing mechanism, agents get trapped in recursive thought loops, failing to recognize that the utility of an optimal decision decays exponentially as time elapses.
  3. 3.
    Epistemic Collapse (Truth): When faced with unresolvable, contradictory evidence, agents lack a rigorous mathematical framework for doubt. Instead of exposing the ambiguity, they average conflicting claims and confidently hallucinate a hybrid synthesis.
In this work, we argue that robust autonomous control requires a formal, decision-theoretic optimization layer that prices information acquisition explicitly. We introduce the Triadic Cognitive Architecture (TCA), which models agent deliberation as a bounded stochastic control problem over a joint belief-routing state space, in which information acquisition is explicitly priced by spatio-temporal and epistemic costs.
Building on related ideas in congestion-aware routing, interval-aware decision-making, and entropic claim resolution digioia2026cascade ; digioia2026learning ; digioia2026entropic , we formally couple the belief state of an agent to the physical friction of its environment. We present a continuous-time formulation for generality (nonlinear filtering and HJB-inspired stopping); in our EMDG implementation we instantiate beliefs on a finite hypothesis set and approximate value-of-information by Monte Carlo rollouts (Section 3.4).
Novelty and contributions.
We contribute: (i) an idealized continuous-time stochastic control formulation in which bounded deliberation is characterized by an HJB-inspired stopping boundary; (ii) a computable discrete instantiation, with formal approximation error bounds (Propositions 1–2) bounding the discrete-to-continuous gap via Itô’s formula and Hoeffding’s inequality, that estimates belief-dependent VOI via rollouts and halts by net utility; (iii) an empirical evaluation on EMDG and a second distinct environment (NSTG) showing that cost-aware selection and principled stopping jointly drive consistent 30+30{+}-point improvements over greedy baselines; and (iv) a continuation-value sweep confirming that $\eta{=}0$ is optimal on EMDG trajectories under high temporal urgency.
Why this is new.
A skeptical reader might object: “This is just rational metareasoning russell1991right (stop when computation cost exceeds utility gain), or classical VOI-based sequential planning, or cost-sensitive bandit stopping with a fancy motivation.” The critical distinction is that these prior frameworks typically treat information-acquisition costs as static scalars fixed independently of the agent’s state. TCA instead makes costs dynamically state-dependent: the spatial friction $\lambda_{S}(C_{t}{+}\Omega(a))$ depends on the live congestion state $C_{t}$, and the temporal friction $\beta\,(t{+}\tau(a))$ is parameterized by the same rate constant $\beta$ that governs the exponential resource decay; both accumulate as deliberation proceeds. TCA jointly models (i) topology- and congestion-dependent information access costs (Space), (ii) time-dependent opportunity costs under delay (Time), and (iii) belief-dependent value-of-information (Truth), within a single stopping framework derived from optimal stopping theory, not ad hoc budget rules. The key contribution is coupling these three state-dependent terms in a unified decision rule and providing a computable rollout-based approximation to the resulting stopping boundary.
The remainder of this paper is structured as follows. Section 2 reviews related work, anchoring TCA within classical bounded rationality and modern agentic AI. Section 3 formalizes the physics of Cognitive Friction and derives the triadic stochastic control objective. Section 4 empirically validates the framework via the Emergency Medical Diagnostic Grid (EMDG). Finally, Section 5 discusses the implications of TCA as a foundational framework for safe, scalable autonomy.
2 Related Work
The Triadic Cognitive Architecture bridges classical theories of bounded computation with modern tool-using autonomous agent design. We briefly contextualize our contributions within these intersecting domains.
2.1 Tool-Using Autonomous Agents
Sequential tool-use and query-then-decide loops have been widely adopted in modern agentic architectures. Frameworks such as ReAct yao2023react , Tree-of-Thoughts yao2023tot , and Reflexion shinn2023reflexion have demonstrated that interleaving reasoning with environmental observations significantly improves task success rates. However, these frameworks rely on arbitrary stopping constraints (e.g., token limits or maximum step counts) wang2023survey , treating each query as free regardless of latency or congestion. Evaluation benchmarks such as AgentBench liu2023agentbench effectively quantify agent failure modes in interactive settings, but implicitly treat deliberation computation as a free resource. TCA argues that safe autonomy demands that computation itself be optimized as a structural cost.
2.2 Bounded Rationality and Meta-Reasoning
The necessity of pricing deliberation traces back to Simon’s foundational concept of bounded rationality simon1955behavioral , which posits that agents must make decisions under the strict limits of available information, cognitive capacity, and time. This was formally extended into meta-reasoning and bounded optimality by Russell and Wefald russell1991right , who argued that an agent should compute only until the cost of computation outweighs the expected increase in decision utility. In deep learning, Adaptive Computation Time (ACT) graves2016adaptive introduced halting mechanisms for recurrent networks based on internal confidence. TCA extends these classical concepts to tool-using autonomous agents, formalizing sequential tool queries as a continuous-time physical process subject to network congestion and temporal decay.
2.3 Active Inference and Stochastic Control
Our framework shares philosophical DNA with Karl Friston’s Free Energy Principle and Active Inference friston2010free , as well as Yann LeCun’s Joint Embedding Predictive Architecture (JEPA). Both theories propose that intelligent agents interact with their environment to minimize surprise (or expected free energy) via a predictive world model. While foundational, exact Bayesian active inference remains largely computationally intractable in practice. TCA offers a principled and computable alternative by providing a closed-loop stochastic control envelope. By synthesizing non-linear filtering kushner1964differential ; bain2009fundamentals with congestion-aware routing and HJB optimal stopping shiryaev2007optimal ; oksendal2003stochastic , TCA bridges stochastic control theory with practical agent systems engineering.
3 Formalizing Cognitive Friction: The Coupled Stochastic Control Objective
Problem setup.
We consider an agent that acquires information via costly tool queries and must decide both which query to perform next and when to stop deliberating and execute a terminal decision. The agent maintains a belief state over hypotheses and a congestion state tracking accumulated network load.
Intuition (for ML practitioners).
At inference time, a tool-using agent faces two coupled questions: which tool should I query next? and when should I stop querying and act? Most agent loops answer both with heuristics (fixed budgets, step limits, or ad hoc confidence thresholds). TCA instead prices deliberation explicitly: each prospective query has (i) a belief-dependent expected benefit (value-of-information (VOI): expected entropy reduction) and (ii) an environment-dependent cost (latency and congestion). The controller chooses the action maximizing net utility and halts when even the best available query has non-positive net value. The continuous-time HJB formulation provides the principled stopping boundary; EMDG instantiates a computable discrete approximation via rollout-based VOI and a myopic stopping rule.
We define the active cognitive state at time $t$ as $s_{t}=(p_{t},C_{t})$, where $p_{t}\in\Delta^{|\Theta|-1}$ is the agent’s belief over a finite hypothesis set $\Theta$, and $C_{t}\in\mathbb{R}_{\geq 0}$ is the scalar network congestion state (accumulated load). Selecting action $u_{t}\in\mathcal{A}$ increments congestion by the action’s load cost:
$\dot{C}_{t}=\Omega(u_{t}),$ (1)
where $\Omega:\mathcal{A}\to\mathbb{R}_{\geq 0}$ is the load-increment function.
3.1 The observation model and belief dynamics
The agent acquires evidence through noisy tool observations. The chosen action $u_{t}\in\mathcal{A}$ selects which channel to query; each tool produces a different signal profile under each hypothesis. The observation process $Y_{t}$ follows an Itô diffusion:
$dY_{t}=h(\theta,u_{t})\,dt+\sigma\,dW_{t},$ (2)
where $h(\theta,u_{t})\in\mathbb{R}^{d}$ is the expected signal of tool $u_{t}$ under hypothesis $\theta$, $W_{t}\in\mathbb{R}^{d}$ is standard Brownian motion, and $\sigma>0$ is a fixed positive scalar so the noise covariance is $\sigma^{2}I_{d}$. Observation quality depends on which tool is selected but not on congestion; congestion enters exclusively through the cost of acquiring observations (Section 3.2). The belief state $p_{t}$ evolves via the Wonham (finite-state Kushner–Stratonovich) filter, driven by the innovation process $d\nu_{t}=dY_{t}-\bar{h}(p_{t},u_{t})\,dt$, where $\bar{h}(p_{t},u_{t})=\int_{\Theta}h(\theta,u_{t})\,p_{t}(\theta)\,d\theta$ is the belief-weighted mean signal of the selected tool.
3.2 The triadic value function
The agent seeks to maximize expected information gain while minimizing spatio-temporal cognitive friction. We define two explicit, state-dependent cost terms:
$\ell_{S}(C_{t},u_{t})$ $:=\lambda_{S}\bigl(C_{t}+\Omega(u_{t})\bigr),$
$\ell_{T}(t)$ $:=\beta\cdot t.$
Here $\lambda_{S}>0$ is a fixed spatial cost weight. The spatial cost $\ell_{S}$ is state-dependent: it grows with the live congestion $C_{t}$ and the load increment $\Omega(u_{t})$ of the chosen action. The temporal cost $\ell_{T}$ reflects the opportunity cost of delay. We use $\beta t$ as a linearly increasing surrogate for time urgency; this aligns directly with the discrete implementation, where the temporal friction term is $\beta(t+\tau(a))$ (a per-action linear penalty). The exponential decay in the resource metric ($e^{-\beta\tau(a)/100}$ per query in EMDG) is a consequence of accumulated linear penalties and does not require an exponential integral formulation. The value function is defined as the supremum over stopping times $T\geq t$:
$V(p,C,t)=\sup_{T\geq t}\,\mathbb{E}\!\left[\,\int_{t}^{T}\!\left(\frac{1}{2\sigma^{2}}\,\mathbb{E}_{p_{s}}\!\left[\bigl\|h(\theta,u_{s})-\bar{h}(p_{s},u_{s})\bigr\|^{2}\right]-\lambda_{S}\bigl(C_{s}+\Omega(u_{s})\bigr)-\beta\,s\right)ds\;\middle|\;p_{t}=p,\;C_{t}=C\,\right],$ (3)
where $\bar{h}(p_{s},u_{s})$ is the belief-weighted mean signal (Section 3.1) and $\sigma$ is the scalar observation noise (Eq. 2); the term $\frac{1}{2\sigma^{2}}\mathbb{E}_{p_{s}}[\|h(\theta,u_{s})-\bar{h}(p_{s},u_{s})\|^{2}]$ equals the expected instantaneous entropy-reduction rate of the Wonham filter. Different tools yield different $h(\theta,u)$ profiles; the $\sup_{u\in\mathcal{A}}$ in the HJB below selects the tool that maximises net reward, trading off information gain against cost. The state $(p_{t},C_{t},t)$ is the minimal sufficient statistic for future net reward: $p_{t}$ determines expected information gain and $C_{t}$ determines the cost trajectory of future queries.
3.3 The HJB variational inequality and optimal stopping
Let $\mathcal{L}^{u}$ denote the (controlled) infinitesimal generator of the Markov state $(p_{t},C_{t})$ under control $u\in\mathcal{A}$, incorporating: (i) the Wonham (finite-state Kushner–Stratonovich) filtering dynamics in $p_{t}$ (see kushner1964differential ; bain2009fundamentals ); and (ii) the deterministic congestion dynamics $\dot{C}_{t}=\Omega(u)$. By dynamic programming, $V(p,C,t)$ satisfies the Hamilton–Jacobi–Bellman (HJB) variational inequality:
$0=\max\left\{\,\partial_{t}V(p,C,t)+\sup_{u\in\mathcal{A}}\!\left(\mathcal{L}^{u}V(p,C,t)+\frac{1}{2\sigma^{2}}\,\mathbb{E}_{p}\!\left[\bigl\|h(\theta,u)-\bar{h}(p,u)\bigr\|^{2}\right]-\lambda_{S}(C+\Omega(u))-\beta\,t\right),\;-V(p,C,t)\,\right\}.$ (4)
In the discrete EMDG instantiation, the running temporal penalty is approximated by an action-dependent one-step cost proportional to latency $\tau(a)$ (Section 3.4).
3.4 Discrete-time instantiation used in EMDG (rollout VOI + myopic stopping)
While the preceding subsections present an idealized continuous-time formulation (filtering and HJB-inspired stopping), the EMDG experiments instantiate a discrete approximation aligned with tool-using agents.
Notation.
We use $\alpha$ for cost-to-utility scaling, $\beta$ for temporal decay/latency penalty, and $\lambda_{S}$ for the spatial cost weight; $\Omega(a)$ and $\tau(a)$ denote action load and latency. Value-of-information (VOI) is defined as belief-dependent expected entropy reduction, i.e., $\widehat{\Delta H}(a\mid b_{t})=\mathbb{E}[H(b_{t})-H(b_{t+1})\mid a]$, estimated by rollouts in EMDG. In the continuous-time formulation $\beta(t)$ is time-varying; in EMDG we approximate it by a constant per-step rate $\beta$ applied to action latency $\tau(a)$. The logged quantity info_gain denotes the realized one-step entropy reduction $H(b_{t})-H(b_{t+1})$; total info_gain telescopes to $H(b_{0})-H(b_{T})$. In the EMDG implementation, the continuous-time running reward $\frac{1}{2\sigma^{2}}\mathbb{E}_{p}[\|h(\theta,u)-\bar{h}(p,u)\|^{2}]$ is replaced by expected entropy reduction, estimated directly via rollout Bayesian updates: each rollout clones the current belief, samples a Dirichlet-likelihood observation, performs the Bayesian update, and records the entropy drop $H(b_{t})-H(b_{t+1})$.
Table 1: Notation summary for the discrete EMDG instantiation.
Symbol Meaning
$b_{t}\in\Delta^{|\Theta|-1}$ Categorical belief over hypotheses at step $t$
$H(b_{t})$ Shannon entropy of belief $b_{t}$ shannon1948
$\widehat{\Delta H}(a\mid b_{t})$ Rollout-estimated VOI (expected entropy reduction)
$\alpha$ Cost-to-utility scaling (cost_scale)
$\beta$ Per-step temporal decay/latency penalty (beta)
$\lambda_{S}$ Spatial cost weight (lambda_s)
$C_{t}$ Current network congestion at step $t$
$\Omega(a)$ Load increment of action $a$ (load_cost)
$\tau(a)$ Latency of action $a$ (time_cost)
$U(a;\,b_{t},t,C_{t})$ Net utility of action $a$ (stop when $\max_{a}U\leq 0$)
info_gain Realized one-step entropy reduction $H(b_{t})-H(b_{t+1})$
We estimate the VOI term $\widehat{\Delta H}(a\mid b_{t})$ via rollouts. The belief state is a categorical distribution $b_{t}\in\Delta^{|\Theta|-1}$ over a finite hypothesis set. Given an action $a$ (a tool query) and an observation sample, we update $b_{t}$ via Bayes’ rule and compute categorical entropy $H(b)=-\sum_{i}b(i)\log\big(b(i)+\varepsilon\big)$.
To estimate belief-dependent value-of-information, we use Monte Carlo rollouts on cloned states:
$\widehat{\Delta H}(a\mid b_{t})=\frac{1}{K}\sum_{k=1}^{K}\Big(H(b_{t})-H(b_{t+1}^{(k)}(a))\Big).$ (5)
This rollout estimator is a Monte Carlo approximation of the mutual information $I(\Theta;Y\mid a,b_{t})$ between hypotheses and observations under the tool’s observation model, a standard information-theoretic quantity that is well-defined for discrete hypothesis sets without requiring a Fisher information matrix. We then score actions by a net-utility objective that prices congestion and latency:
$U(a;b_{t},t,C_{t})=\widehat{\Delta H}(a\mid b_{t})-\alpha\Big(\lambda_{S}(C_{t}+\Omega(a))+\beta\,(t+\tau(a))\Big),$ (6)
and stop when the best continuation utility is non-positive:
$\texttt{STOP}\;\Longleftrightarrow\;\max_{a\in\mathcal{A}}U(a;b_{t},t,C_{t})\leq 0.$ (7)
Note that the triadic controller optimizes a proxy objective: expected entropy reduction minus spatio-temporal friction. The physical resource metrics reported in our experiments (patient viability and system integrity) are external environmental variables that decay with elapsed time. By explicitly pricing delay through the temporal term $\beta\,(t+\tau(a))$, the controller is designed to trade off information gain against preservation of these external metrics without requiring them to appear directly in the agent’s internal belief state. This myopic stopping rule omits continuation value (equivalently, setting a one-step lookahead weight $\eta=0$ in a one-step lookahead formulation); we empirically validate the optimality of this choice in Section 4.8. Here, $C_{t}$ denotes the congestion state (Eq. 1), while $\eta$ is a separate scalar lookahead weight (set to zero in our experiments).
On the discretization gap.
The idealized HJB variational inequality optimizes a continuation value over future information paths. Our EMDG controller uses a myopic rule ($\eta=0$), i.e., it compares immediate VOI against immediate spatio-temporal cost and halts when the best net utility is non-positive. This choice is deliberate: it yields a simple, inference-time controller that is robust and computationally inexpensive. In principle, continuation value can be approximated by multi-step rollouts or learned value functions; we empirically test $\eta\in\{0.0,0.1,0.3,0.5\}$ in Section 4.8 and find that continuation value is identically zero throughout the EMDG trajectory, empirically demonstrating that $\eta=0$ is optimal in this environment.
Proposition 1 (Rollout–HJB approximation error).
To formalize the bridge from continuous-time HJB rewards to the discrete rollout estimator, we bound the approximation error in two parts. The rollout VOI estimator $\widehat{\Delta H}(a\mid b_{t})$ (Eq. 5) approximates the filtered information-rate term $\frac{\tau(a)}{2\sigma^{2}}\,\mathbb{E}_{p}\bigl[\|h(\theta,a)-\bar{h}(p,a)\|^{2}\bigr]$ (the leading-order continuous-time running reward in Section 3.2) with a two-part error bound.
  1. (i)
    Monte Carlo error. By the law of large numbers, $\widehat{\Delta H}(a\mid b_{t})$ is an unbiased estimator of the discrete mutual information $I(\Theta;Y\mid a,b_{t})=\mathbb{E}[H(b_{t})-H(b_{t+1}(a))]$. Since $|H(b_{t})-H(b_{t+1})|\leq\log|\Theta|$ (bounded range), Hoeffding’s inequality gives
    $\bigl|\widehat{\Delta H}(a\mid b_{t})-I(\Theta;Y\mid a,b_{t})\bigr|=O(K^{-1/2})$
    with probability at least $1-2\exp(-K\varepsilon^{2}/(2\log^{2}|\Theta|))$ for any $\varepsilon>0$.
  2. (ii)
    Temporal discretization error. By Itô’s formula applied to $H(p_{t})$ under the Wonham filter bain2009fundamentals , the infinitesimal drift of entropy is:
    $dH(p_{t})=-\frac{1}{2\sigma^{2}}\,\mathbb{E}_{p_{t}}\bigl[\|h(\theta,u_{t})-\bar{h}(p_{t},u_{t})\|^{2}\bigr]\,dt+dM_{t},$
    where $M_{t}$ is a zero-mean martingale. The drift term is the signal-variance under the current belief ($\sigma^{-2}$ scaled), which matches the leading-order continuous-time running reward in Section 3.2. Integrating over $t,\,t{+}\tau(a)]$ and taking expectations (the martingale vanishes):
    $I(\Theta;Y\mid a,b_{t})\;=\;\frac{\tau(a)}{2\sigma^{2}}\,\mathbb{E}_{p}\bigl[\|h(\theta,a)-\bar{h}(p,a)\|^{2}\bigr]+O\!\bigl(\tau(a)^{2}\bigr),$
    where the residual reflects the change in $\Sigma_{p_{t}}$ over the action step. Under standard regularity assumptions for finite-state filtering (bounded $h$, $\sigma>0$, and local Lipschitz continuity of $\mathbb{E}_{p}[\|h(\cdot,u)-\bar{h}(p,u)\|^{2}]$ in $p$ along the Wonham dynamics), there exists a constant $L_{\Sigma}\geq 0$, depending on bounds on $h$ and on the rate of change of the signal variance under the chosen action, such that
    $\biggl|I(\Theta;Y\mid a,b_{t})-\frac{\tau(a)}{2\sigma^{2}}\,\mathbb{E}_{p}\bigl[\|h(\theta,a)-\bar{h}(p,a)\|^{2}\bigr]\biggr|\leq L_{\Sigma}\,\tau(a)^{2}.$
Combining (i) and (ii), the aggregate error of the rollout estimator with respect to the HJB information-rate term is bounded by $L_{\Sigma}\,\tau(a)^{2}+O(K^{-1/2})$. (The spatio-temporal cost terms are computed exactly; only the information-rate term is approximated.) The bound is tightest for short-latency actions (e.g., Hematology Lab, $\tau{=}5$) and degrades for long-latency actions (e.g., MRI, $\tau{=}45$); in practice, TCA predominantly selects low-$\tau$ actions (Table 2), keeping execution within the well-approximated regime.
On model idealization.
The continuous-time formulation provides an idealized optimal-stopping envelope for bounded deliberation under state-dependent costs. Our implemented controller is a discrete-time approximation that estimates VOI by Monte Carlo belief updates and applies a net-utility stopping rule. The two formulations are connected by a consistency limit: as $K\to\infty$ the VOI estimate converges to mutual information, and for short action latencies ($\tau(a)\to 0$) the discrete information gain matches the leading-order continuous-time reward term (Proposition 1). We therefore interpret the controller’s myopic net-utility test as a first-order approximation of the HJB stopping boundary rather than solving the full HJB value function.
Proposition 2 (First-order approximation to the HJB stopping boundary).
Let $V^{\eta}(p,C,t)$ denote the one-step lookahead value under continuation weight $\eta\geq 0$. At $\eta=0$, the HJB optimality condition $V^{0}(p,C,t)\leq 0$ reduces to
$\max_{a\in\mathcal{A}}\,U(a;\,b_{t},t,C_{t})\leq 0,$
i.e., the myopic net-utility rule (Eq. 7) corresponds to the $\eta=0$ (one-step) approximation to the HJB stopping boundary. Correction terms of order $O(\eta)$ correspond to discounted future utility; Section 4.8 verifies empirically that these are zero throughout EMDG trajectories, consistent with approximate dynamic programming bertsekas2012approximate .
Theorem 1 (Optimal stopping in the idealized TCA model).
Assume (i) $h:\Theta\times\mathcal{A}\to\mathbb{R}^{d}$ is bounded, $\sigma>0$, $\Omega:\mathcal{A}\to\mathbb{R}_{\geq 0}$ is bounded, and $\lambda_{S}>0$, $\beta>0$; and (ii) admissible controls are adapted to the observation filtration, with stopping times taken with respect to the same filtration. Then the idealized continuous-time triadic control problem admits an optimal stopping time. Moreover, letting $D=\{(p,C,t):V(p,C,t)=0\}$ denote the stopping region, the minimal optimal stopping time is the first hitting time
$T^{*}=\inf\{s\geq t:(p_{s},C_{s},s)\in D\}.$
Corollary 1 (Monotone stopping region).
Under the hypotheses of Theorem 1, the stopping region $D$ is an upper set in $(C,t)$:
$(p,C,t)\in D\;\Longrightarrow\;(p,C^{\prime},t^{\prime})\in D\quad\text{for all }C^{\prime}\geq C,\;\;t^{\prime}\geq t.$
Equivalently, for each $p\in\Delta^{|\Theta|-1}$ there exist free-boundary functions $\bar{C}(p,t)\in[0,\infty]$ and $\bar{t}(p,C)\in[0,\infty]$ such that continuation is optimal if and only if $C_{t}<\bar{C}(p_{t},t)$ and $t<\bar{t}(p_{t},C_{t})$.
Proof sketch (Corollary 1).
(Monotone in $C$). Since $\dot{C}_{t}=\Omega(u_{t})\geq 0$ is independent of $C$, shifting the initial congestion from $C$ to $C^{\prime}=C+\delta$ ($\delta>0$) translates every future $C_{s}$ by $\delta$, raising the spatial cost $\lambda_{S}(C_{s}{+}\Omega(u_{s}))$ by $\lambda_{S}\delta>0$ at every instant. For all admissible stopping times $T\geq t$ and controls:
$\mathbb{E}\!\left[\int_{t}^{T}\!\bigl(r_{s}-\lambda_{S}(C_{s}{+}\delta{+}\Omega(u_{s}))-\beta s\bigr)ds\right]\;\leq\;\mathbb{E}\!\left[\int_{t}^{T}\!\bigl(r_{s}-\lambda_{S}(C_{s}{+}\Omega(u_{s}))-\beta s\bigr)ds\right].$
Taking the supremum over $T$: $V(p,C^{\prime},t)\leq V(p,C,t)=0$. Since $V\geq 0$ (immediate stop yields zero value), $(p,C^{\prime},t)\in D$.
(Monotone in $t$). The Wonham filter and congestion dynamics are time-homogeneous ($h(\theta,u)$ and $\Omega(u)$ carry no explicit calendar-time dependence). Starting at $t^{\prime}=t+\delta$ with state $(p,C)$ is therefore equivalent to starting at $t$ with an extra $\beta\delta$ cost per unit time:
$V(p,C,t+\delta)=\sup_{T^{\prime}\geq t}\,\mathbb{E}\!\left[\int_{t}^{T^{\prime}}\!\bigl(r_{s}-\lambda_{S}(C_{s}{+}\Omega(u_{s}))-\beta s-\beta\delta\bigr)ds\right]\;\leq\;V(p,C,t).$
[…]
Operational interpretation.
Since $C_{t}$ is non-decreasing and $t$ advances monotonically, Corollary 1 implies the stopping boundary is absorbing: once the state enters $D$, no future trajectory can re-enter the continuation region. This one-way ratchet is a structural consequence of TCA’s cost formulation, absent from classical meta-reasoning frameworks russell1991right where stop/continue decisions may cycle under non-monotone utility estimates. Corollary 1 also provides theoretical grounding for the myopic stopping rule (Eq. 7): once the best available action yields non-positive net utility, this condition is guaranteed to persist.
Remark (theory vs. instantiation).
Theorem 1 and Corollary 1 concern the idealized continuous-time formulation. Our EMDG experiments implement a discrete approximation (Section 3.4) via rollout-based VOI estimation and a myopic stopping rule (Propositions 1–2).
4 Empirical Validation: The Emergency Medical Diagnostic Grid (EMDG)
To demonstrate the necessity of constrained spatio-temporal and epistemic bounds, we evaluate TCA in the Emergency Medical Diagnostic Grid (EMDG). Rather than a high-fidelity real-world benchmark, EMDG is designed as an idealized, synthetic testbed to strictly isolate spatial, temporal, and epistemic friction dynamics under controlled, reproducible conditions. In this simulation, an autonomous diagnostic agent is tasked with identifying a critical pathology and prescribing an intervention.
The environment imposes strict physical realities:
  • Epistemic ambiguity. The agent begins with a uniform hypothesis prior over five highly fatal pathologies ($B_{0}$).
  • Spatial topology. The agent retrieves evidence by routing queries to distributed hospital sub-systems (e.g., Hematology Lab, MRI network, Patient History). Each sub-system has a distinct structural load cost ($\Omega$).
  • Temporal decay. Patient survivability decays exponentially with elapsed time. Some queries (e.g., rapid blood tests) cost 5 time-steps; others (e.g., MRI) cost 45 time-steps.
00.50.511Entropy HHEMDG Monte Carlo (N=200): mean trajectoriesReAct (Greedy)Triadic Control00.50.5111.51.5222.52.5333.53.54460608080100100StepPatient viability
Figure 1: EMDG Monte Carlo (N=200): mean entropy and patient viability trajectories over a fixed horizon. Step corresponds to discrete tool-query iterations; trajectories are padded after stopping by carrying forward the terminal state, ensuring step-wise consistency (e.g., time cannot decrease).
We evaluate a standard greedy tool-use baseline (utilizing a ReAct-style Thought $\to$ Act $\to$ Observe loop yao2023react ) against the Triadic Cognitive Architecture (TCA).
Evaluation integrity.
To prevent estimator/trajectory coupling, we decouple randomness by using a per-environment RNG for the realized transition and observation draws, and a separate independent RNG stream for value-of-information (VOI) rollouts on cloned states. While HJB theory motivates the stopping boundary, our EMDG controller instantiates a rollout-based approximation of belief-dependent VOI with a net-utility halting condition.
4.1 The failure of the unconstrained agent (ReAct)
Operating without cognitive friction, the baseline ReAct agent greedily maximizes expected entropy reduction and does not price congestion or latency into action selection. At step 0 it initiates a query to the MRI network.
Spatial failure.
Because it ignores the load and congestion penalties associated with routing, it is willing to select high-load tools whenever their immediate estimated gain is maximal.
Temporal failure.
Because MRI incurs a 45 time-step latency, viability decays exponentially with elapsed time before a low-latency intervention would be taken.
Epistemic failure.
The baseline lacks a stopping rule that compares marginal entropy reduction to time-sensitive viability loss; as a result, it continues to query even when entropy is already low and additional deliberation is not worth the delay.
4.2 The success of the triadic agent (TCA)
The TCA dynamically evaluates the unified objective function at $t=0$. The MRI offers high expected entropy reduction (VOI), but its high temporal cost (latency $\tau(a)$) and spatial congestion increment $\Omega(a)$ yield negative net cognitive utility.
Action mix.
At step 0, greedy ReAct selects MRI_Network in 100% of seeds, whereas the triadic controller selects Hematology_Lab in 100% of seeds, reflecting a consistent preference for low-latency evidence acquisition under cognitive friction.
Instead, TCA routes a lightweight query to the Hematology Lab ($\tau=5$, low $\Omega$). Over a small number of low-latency queries, it reaches a regime where the marginal expected entropy reduction (VOI) is eclipsed by spatio-temporal cost, and the policy halts.
Concretely, TCA estimates VOI by rollout entropy reduction (Eq. 5), scores queries by net utility (Eq. 6), and halts when the best continuation utility is non-positive (Eq. 7):
$U(a;b_{t},t,C_{t})=\widehat{\Delta H}(a\mid b_{t})-\alpha\!\left(\underbrace{\lambda_{S}(C_{t}+\Omega(a))}_{\text{spatial friction}}+\underbrace{\beta\,(t+\tau(a))}_{\text{temporal friction}}\right).$
We report mean $\pm$ 95% confidence intervals computed as $1.96\,\hat{\sigma}/\sqrt{N}$ over $N=200$ seeds, using terminal-per-seed values (last logged row per seed).
Table 2: EMDG results (N=200). Terminal values are taken from the last logged row per seed. Total information gain is the per-episode sum of info_gain; in this setup it equals entropy reduction $H(b_{0})-H(b_{T})$ (a telescoping sum). At step 0, ReAct selects MRI_Network in 100% of seeds while TCA selects Hematology_Lab in 100% of seeds. Bold marks the best value for the two primary objectives (Time and Viability); lower entropy and higher information gain in ReAct reflect expensive over-querying, not better performance.
Agent Time Viability Entropy Accuracy $p_{\mathrm{true}}$ Total information gain
ReAct (Greedy) 114.5 $\pm$ 3.1 56.76 $\pm$ 0.89 0.0386 $\pm$ 0.0038 1.00 $\pm$ 0.00 0.9941 $\pm$ 0.0007 1.5708 $\pm$ 0.0038
Triadic Control 14.5 $\pm$ 0.4 93.03 $\pm$ 0.19 0.1714 $\pm$ 0.0119 1.00 $\pm$ 0.00 0.9643 $\pm$ 0.0032 1.4380 $\pm$ 0.0119
Note.
In EMDG, info_gain is logged per step as $H(b_{t})-H(b_{t+1})$, so per-episode total information gain telescopes to $H(b_{0})-H(b_{T})$. Because $b_{0}$ is fixed by the uniform prior, the dispersion (and thus CI width) of total information gain matches that of terminal entropy.
Scalability considerations.
The EMDG environment is intentionally small to isolate bounded-deliberation effects under congestion and latency. Accuracy is intentionally saturated: the task is constructed such that correct inference is achievable with sufficient queries. This isolates the decision problem (when and how to acquire information) from the orthogonal problem of model capability under ambiguous evidence, allowing us to evaluate cost-efficient stopping as a pure decision-theoretic quantity. Differences in viability and deliberation time are therefore attributable solely to stopping quality and action selection, not to inference accuracy. Appendix B further evaluates TCA on larger, randomized action spaces ($|\mathcal{A}|\in\{5,10,20\}$). In larger tool ecosystems, naive Monte Carlo VOI rollouts can be costly. However, VOI estimation can be amortized via (i) learned surrogates for expected entropy reduction, (ii) caching/reuse of rollouts for similar belief states, (iii) restricting candidate tools via cheap screening heuristics, or (iv) using analytic approximations (e.g., Fisher-information or local linearization) when available. These approximations align with the TCA view: the controller requires priced estimates of marginal information value, not necessarily exact computation. In real LLM deployments, belief $b_{t}$ can be approximated from model output logits or next-token probabilities over hypothesis labels; VOI can then be estimated via LLM ensemble disagreement or by sampling tool-conditioned responses. Cost terms ($\tau(a)$, $\Omega(a)$) map directly to measured API latency and token-consumption statistics, grounding TCA in observable runtime quantities.
Generalization to heterogeneous tool ecosystems.
We evaluate TCA under a second, structurally distinct environment in Section 4.7 and confirm consistent behavior across different domain, tool profiles, and urgency regime.
4.3 Hyperparameters and reproducibility
In EMDG, the cost-to-utility scaling $\alpha$, temporal decay $\beta$, and spatial cost weight $\lambda_{S}$ are fixed across all seeds to reflect a time-sensitive diagnosis regime under congestion. In our current implementation these correspond to cost_scale ($\alpha=0.01$), beta ($\beta=0.5$), and lambda_s ($\lambda_{S}=0.8$), respectively. Section 4.6 reports a univariate sensitivity sweep over all three parameters.
4.4 Ablation study
To isolate the contribution of each component of the triadic controller, we run a controlled ablation sweep over $N=200$ Monte Carlo seeds. We ablate (i) the HJB stopping rule by disallowing STOP (No-Stop), (ii) the spatial term by setting the spatial friction weight to zero (No-Space), (iii) the temporal term by setting the temporal friction weight to zero (No-Time), and (iv) the congestion feedback term by ignoring the live congestion state in the spatial friction calculation (No-Congestion). Figure 2 reports terminal patient viability and terminal deliberation time (means $\pm 1\sigma$).
[…]
Figure 2: EMDG ablations (N=200): terminal patient viability and terminal deliberation time across controller variants. Bars show means and error bars denote $\pm 1\sigma$ over seeds. The full triadic controller achieves the highest viability while minimizing deliberation time, illustrating the necessity of jointly pricing space, time, and enforcing an explicit stopping boundary.
4.5 Purposive stopping baselines
The ablations in Section 4.4 remove individual TCA cost terms while keeping TCA’s joint selection policy intact. To answer the distinct question of whether TCA’s advantage is explained solely by its stopping criterion, independent of cost-aware selection, we evaluate two purposive baselines, each pairing the same greedy selection policy as ReAct (always query the highest-expected_gain tool, MRI_Network) with a principled stopping rule calibrated to match TCA’s behavior:
  • Entropy-Threshold (EnT, $\tau_{H}{=}0.17$). Stop before querying when $H(b_{t})<\tau_{H}$, where $\tau_{H}$ is TCA’s mean terminal entropy ($0.1714$). Tests whether stopping at the right uncertainty level is sufficient.
  • Fixed-K (K=3K{=}3). Stop after exactly $K$ queries, where $K$ equals TCA’s mean query count before Stop. Tests whether matching TCA’s query budget alone explains the gain.
Both baselines are matched to TCA in one dimension of its stopping behavior while using cost-blind (greedy) query selection. Table 3 reports results ($N=200$).
Table 3: Extended comparison: TCA vs. purposive stopping baselines (N=200). Entropy-Threshold uses $\tau_{H}{=}0.17$ (TCA mean terminal entropy); Fixed-K uses K=3K{=}3 (TCA mean query count before Stop). Both baselines use greedy (cost-blind) query selection identical to ReAct. TCA values reproduced from Table 2 for convenience. Bold marks the best per metric.
Agent Time Viability Entropy Accuracy
ReAct (Greedy) 114.5 $\pm$ 3.1 56.76 $\pm$ 0.89 0.0386 $\pm$ 0.0038 1.00 $\pm$ 0.00
Entropy-Threshold ($\tau_{H}{=}0.17$) 102.6 $\pm$ 2.8 60.17 $\pm$ 0.80 0.0687 $\pm$ 0.0060 1.00 $\pm$ 0.00
Fixed-K (K=3K{=}3) 135.0 $\pm$ 0.0 50.92 $\pm$ 0.00 0.0152 $\pm$ 0.0021 1.00 $\pm$ 0.00
Triadic Control (TCA) 14.5 $\pm$ 0.4 93.03 $\pm$ 0.19 0.1714 $\pm$ 0.0119 1.00 $\pm$ 0.00
Interpretation.
Purposive stopping alone does not close the viability gap. EnT reduces mean deliberation time from 114.5 to 102.6, yet viability improves only marginally ($56.8\to 60.2$), still 33 points below TCA. Fixed-K performs worse than unconstrained ReAct ($50.9$ vs. $56.8$) because forcing three MRI queries exhausts the time budget without the patient receiving timely treatment. TCA’s 33-point advantage over EnT and 42-point advantage over Fixed-K arise from its cost-aware query selection: by pricing latency and congestion into every action-utility estimate, the triadic controller routes to Hematology_Lab on every seed at step 0, whereas all greedy baselines select MRI_Network. These results indicate that selection and stopping must be optimized jointly; stopping alone is insufficient under high-latency tool profiles.
4.6 Parameter sensitivity
Figure 3 reports terminal patient viability under univariate sweeps of each TCA hyperparameter, with the remaining two held at their defaults ($N=200$ seeds per setting).
Cost-scale ($\alpha$). Higher $\alpha$ increases the penalty on spatio-temporal cost, causing TCA to halt earlier (mean time $44.3\to 11.5$) and lifting viability ($80.3\to 94.4$, monotone). This confirms that $\alpha$ directly controls the viability–deliberation tradeoff: under-penalising cost leads the agent to over-query expensive tools, while the default ($\alpha=0.01$) and higher values consistently avoid the high-latency MRI path.
Temporal decay ($\beta$). The effect of $\beta$ is non-monotone because it governs both the environment’s physical viability decay and the temporal friction term simultaneously. At $\beta=0.25$ the patient degrades slowly, yielding high viability (96.3) even with moderate deliberation time (15.0). At $\beta=1.0$ the rapid physical decay dominates: despite TCA stopping earlier (time 13.2), every time-step costs more, pushing viability to 87.7. TCA remains well above ReAct ($56.8$) across the entire range.
Spatial weight ($\lambda_{S}$). Higher $\lambda_{S}$ penalises congested access more heavily, shortening deliberation ($22.2\to 13.3$) and improving viability ($89.7\to 93.6$). The improvement plateaus between $\lambda_{S}=0.80$ and $1.20$, indicating that once the spatial term is large enough to reliably select low-congestion tools, further increases yield diminishing returns.
Across all 7 configurations, accuracy remains 1.0 and TCA always outperforms ReAct on viability, confirming robustness over the tested ranges. The results support $\alpha$ as the primary tuning knob for the viability–deliberation tradeoff, with $\beta$ and $\lambda_{S}$ best set to reflect the physical environment’s temporal urgency and congestion profile.
[…]
Figure 3: Univariate parameter sensitivity: terminal patient viability (mean $\pm$ 95% CI, N=200N{=}200 seeds per setting). Each panel varies one parameter while holding the others at default ($\alpha{=}0.01$, $\beta{=}0.50$, $\lambda_{S}{=}0.80$); the middle bar in each panel corresponds to the default. Dashed line in the $\alpha$ panel marks ReAct viability (56.8) for reference. Accuracy remains 1.0 across all 7 configurations.
4.7 Cross-environment generalization: NSTG
To assess whether TCA’s advantage is specific to the medical emergency domain, we evaluate an independent second environment: the Network Security Triage Grid (NSTG). NSTG models a SOC agent identifying a network threat type (5 categories: Ransomware, APT, DataExfiltration, DDoS Amplification, InsiderThreat) and issuing containment. The resource metric is system_integrity $\in[0,100]$, decaying exponentially with elapsed time according to the same formula as EMDG patient viability. The tool suite is structurally distinct: QuickScan ($\tau{=}4$, $\Omega{=}3$, expected gain $0.40$) versus FullForensics ($\tau{=}60$, $\Omega{=}70$, expected gain $1.30$). The cost regime also differs: $\beta{=}0.30$ (lower temporal urgency than EMDG’s $0.50$, reflecting that incident response is less time-critical than emergency medicine), $\lambda_{S}{=}0.90$, $\alpha{=}0.015$. A congestion shock ($\times 3$ at step 2) is applied, simulating a burst of competing requests from other agents in the network.
TCA selects QuickScan in 100% of seeds at step 0, whereas ReAct selects FullForensics in 100% of seeds; this mirrors the EMDG latency-driven failure mode.
Table 4 reports results (N=200N{=}200). TCA achieves $97.18{\pm}0.08$ system integrity vs. $64.08{\pm}0.80$ for ReAct (a 33.1-point gap), and halts $\approx 16\times$ faster (time $9.5{\pm}0.3$ vs. $149.7{\pm}4.2$). Accuracy is $1.00$ for both agents. The slightly smaller gap than EMDG (33.1 vs. 36.3 points) is expected: lower $\beta$ means each query costs less temporal friction, so even ReAct’s high-latency path loses less integrity per time-step. The advantage direction and relative magnitude are fully consistent across both domains, supporting the hypothesis that cognitive friction pricing constitutes a domain-independent principle.
Table 4: NSTG results (N=200). Resource metric is system integrity (same exponential decay formula as EMDG patient viability, with $\beta{=}0.30$). At step 0, ReAct selects FullForensics in 100% of seeds while TCA selects QuickScan in 100% of seeds. Bold marks the best per metric.
Agent Time Integrity Entropy Accuracy $p_{\mathrm{true}}$ Total information gain
ReAct (Greedy) 149.7 $\pm$ 4.2 64.08 $\pm$ 0.80 0.0376 $\pm$ 0.0037 1.00 $\pm$ 0.00 0.9943 $\pm$ 0.0006 1.5718 $\pm$ 0.0037
Triadic Control 9.5 $\pm$ 0.3 97.18 $\pm$ 0.08 0.3230 $\pm$ 0.0213 1.00 $\pm$ 0.00 0.9180 $\pm$ 0.0083 1.2864 $\pm$ 0.0213
Table 5 provides a side-by-side summary across both environments.
Table 5: Cross-environment summary: TCA vs. ReAct (Greedy), N=200N{=}200 seeds per environment. Resource metric == patient viability (EMDG) / system integrity (NSTG); both decay exponentially with elapsed time. Bold marks the best per row.
Environment Agent Time Resource Entropy Accuracy
EMDG ($\beta{=}0.50$) ReAct (Greedy) 114.5 $\pm$ 3.1 56.76 $\pm$ 0.89 0.0386 $\pm$ 0.0038 1.00
Triadic Control 14.5 $\pm$ 0.4 93.03 $\pm$ 0.19 0.1714 $\pm$ 0.0119 1.00
NSTG ($\beta{=}0.30$) ReAct (Greedy) 149.7 $\pm$ 4.2 64.08 $\pm$ 0.80 0.0376 $\pm$ 0.0037 1.00
Triadic Control 9.5 $\pm$ 0.3 97.18 $\pm$ 0.08 0.3230 $\pm$ 0.0213 1.00
4.8 One-step continuation value ($\eta>0$)
Section 3.4 discloses that the EMDG controller uses $\eta{=}0$ (myopic stopping) as a deliberate design choice. We now close this gap empirically by sweeping $\eta\in\{0.0,\,0.1,\,0.3,\,0.5\}$ on EMDG (N=200N{=}200 seeds), implementing one-step lookahead via the belief-cloning rollout.
All four configurations produce identical terminal statistics: time $14.4{-}14.5{\pm}0.4$, viability $93.02{-}93.04{\pm}0.20$, entropy $0.170{-}0.172{\pm}0.012$, accuracy $1.00$. This is not numerical coincidence: rollout inspection confirms that $\texttt{continuation\_value}(a){=}0$ for all actions at every step. In EMDG, after each Hematology_Lab query the agent is already at or past the myopic stopping boundary: the best available next action has non-positive expected net utility, so the $\eta$-weighted continuation term is zero and does not affect action scores or stopping decisions.
This provides a stronger-than-expected result: the myopic rule is not merely near-optimal but exactly optimal in EMDG for any $\eta\in[0,1]$. Intuitively, in high-urgency regimes (large $\beta$), the temporal penalty forces the stopping boundary to coincide with the one-step-ahead boundary; the zero-continuation property reflects this regime. For lower-urgency environments (e.g., NSTG with $\beta{=}0.30$), continuation value may be non-zero; we leave systematic investigation of multi-step lookahead in moderate-urgency settings to future work.
4.9 Congestion decay robustness ($\kappa>0$)
The Limitations section notes that the default EMDG model uses monotone congestion accumulation ($\dot{C}_{t}=\Omega(u_{t})$, $\kappa{=}0$). Real networks exhibit congestion relief between queries (queue drain, TCP backoff). We close this gap by extending the congestion dynamics to the discrete Ornstein–Uhlenbeck model
$C_{t+\tau(a)}=C_{t}\,e^{-\kappa\,\tau(a)}+\Omega(a),$
where $\kappa\geq 0$ is the drain rate ($\kappa{=}0$ recovers the original model exactly). We sweep $\kappa\in\{0.00,\,0.05,\,0.10\}$ over N=200N{=}200 seeds with all other parameters at their defaults.
Table 6 reports terminal statistics. ReAct is unaffected by $\kappa$ (it ignores congestion costs by construction and always selects MRI_Network). TCA viability remains stable at 92.792.793.093.0 across the sweep (a <0.4<0.4-point variation well within the 95% CI), and the ${\approx}36$-point advantage over ReAct is fully preserved. TCA deliberation time increases by ${\approx}0.7$ steps at $\kappa{=}0.10$: reduced congestion lowers spatial cost, occasionally making one additional Hematology_Lab query worthwhile before the stopping boundary is hit. This is the expected and desirable behaviour of a cost-aware controller.
Table 6: Congestion-decay robustness: TCA vs. ReAct (Greedy) for $\kappa\in\{0.00,0.05,0.10\}$ (N=200N{=}200 seeds). $\kappa{=}0.00$ replicates Table 2 exactly. Bold marks best viability per $\kappa$ level.
$\kappa$ Agent Time Viability Accuracy
$0.00$ ReAct (Greedy) $114.5\pm 3.1$ $56.76\pm 0.89$ $1.00$
$0.00$ Triadic Control $\mathbf{14.5\pm 0.4}$ $\mathbf{93.03\pm 0.19}$ $1.00$
$0.05$ ReAct (Greedy) $114.5\pm 3.1$ $56.76\pm 0.89$ $1.00$
$0.05$ Triadic Control $\mathbf{14.8\pm 0.5}$ $\mathbf{92.89\pm 0.21}$ $1.00$
$0.10$ ReAct (Greedy) $114.5\pm 3.1$ $56.76\pm 0.89$ $1.00$
$0.10$ Triadic Control $\mathbf{15.2\pm 0.5}$ $\mathbf{92.72\pm 0.22}$ $1.00$
5 Implications: TCA as a Foundational Framework for Bounded Agent Reasoning
The Triadic Cognitive Architecture (TCA) provides a blueprint for transitioning from disembodied pattern matchers to grounded, autonomous agents. By formalizing cognitive friction, TCA introduces a mathematically rigorous safety mechanism against the runaway resource consumption inherent in unconstrained LLM reasoning loops.
5.1 Coupling Space, Time, and Epistemic Truth
Within the TCA framework, an agent’s reasoning is strictly bounded by three interacting modules (see Figure 4):
  • Space (Topological Grounding): Tools are not treated as a flat list of interchangeable APIs but as resources with distinct latency and congestion costs that depend on the agent’s current network state. In both the continuous-time formulation (Section 3) and the EMDG/NSTG implementations, the spatial friction term $\ell_{S}(C_{t},a)=\lambda_{S}(C_{t}{+}\Omega(a))$ prices access by live congestion state $C_{t}$ (accumulated network load, updated after each query) and structural load increment $\Omega(a)$. This prevents API spamming and protects fragile sub-agents from cascade failures.
  • Time (Temporal Consistency): Infinite context lengths do not circumvent the reality that the physical world evolves while the agent deliberates. The exponential decay factor $\beta(t)$ ensures that the agent respects the temporal opportunity cost of computation, intrinsically preventing infinite deliberation loops.
  • Truth (Epistemic Maintenance): The agent maintains a calibrated categorical belief $b_{t}$ over hypotheses, updated via Bayes’ rule after each tool observation. In the idealized formulation (Section 3), the running reward is $\frac{1}{2\sigma^{2}}\mathbb{E}_{p}[\|h(\theta,u)-\bar{h}(p,u)\|^{2}]$ (tool-dependent, since different tools probe different signal channels); in EMDG/NSTG, this is instantiated as rollout-estimated entropy reduction $\widehat{\Delta H}(a\mid b_{t})$. The stopping rule halts deliberation once marginal entropy reduction no longer justifies the spatio-temporal cost, naturally avoiding the trap of averaging contradictory evidence at the cost of the resource metric.
Taken together, these three forms of friction (spatial, temporal, and epistemic) define a bounded control envelope for autonomous reasoning. Rather than optimizing accuracy in isolation, TCA determines when to act, what to query, and how long to deliberate by explicitly trading off expected entropy reduction against physical and temporal cost. This triadic coupling provides a principled mechanism for preventing over-deliberation, network saturation, and epistemic overconfidence in autonomous agents.
Space(state)Time(dynamics)EpistemicTruthperceptionpredictionrevision
Figure 4: Triadic Cognitive Architecture (TCA): a coupled loop between Space (state), Time (dynamics), and Epistemic Truth (belief revision).
5.2 The Inference-Time Controller
The core of TCA is an inference-time controller motivated by HJB optimal stopping. It continuously aggregates the spatial routing cost and temporal decay, dynamically weighing them against expected entropy reduction (VOI). This controller acts as an autonomic nervous system for the AI: just as biological organisms intuitively halt exploration when caloric costs exceed potential rewards, the TCA agent halts API queries and triggers execution once the marginal value of additional evidence is eclipsed by spatio-temporal cost.
5.3 Positioning Relative to Cognitive Theories
Extensive work has been done in prompt-driven multi-step reasoning, notably ReAct yao2023react , Tree-of-Thoughts yao2023tot , and Reflexion shinn2023reflexion . However, these frameworks rely on heuristic, discrete hypothesis sets, ignore network-dependent physical costs, and lack principled mathematical criteria for optimal stopping liu2023agentbench ; wang2023survey .
The TCA framework shares philosophical DNA with Karl Friston’s Free Energy Principle and Active Inference friston2010free , as well as Yann LeCun’s Joint Embedding Predictive Architecture (JEPA). Both propose that intelligent agents seek to minimize surprise in a predictive world model. However, Active Inference relies on variational bounds that remain largely intractable in practice. TCA offers a principled and computable alternative by providing a closed-loop stochastic control envelope. By synthesizing non-linear filtering kushner1964differential ; bain2009fundamentals with congestion-aware routing and HJB optimal stopping shiryaev2007optimal ; oksendal2003stochastic , TCA bridges stochastic control theory with practical agent systems engineering.
5.4 Limitations
We list the main gaps between the idealized framework (Section 3) and the current empirical instantiation.
Small action space. EMDG and NSTG each offer two non-trivial tools plus Stop. The key experimental contrast (latency-aware vs. greedy selection) is cleanly isolated in this setting. Appendix B addresses this gap empirically: a synthetic scaling sweep over $|\mathcal{A}|\in\{5,10,20\}$ shows TCA selecting non-minimum-latency tools in 76–90% of configurations and consistently outperforming ReAct (viability gap $\geq+9.7$ points) across all action-space sizes tested. VOI estimation via K=32K{=}32 rollouts is $O(K{\times}|\mathcal{A}|)$; in a system with hundreds of tools, surrogate models or screening heuristics (Section 4.2) would be required.
Monotone congestion accumulator. The default model ($\kappa{=}0$) accumulates congestion monotonically. Section 4.9 extends this to a discrete Ornstein–Uhlenbeck drain model ($\kappa{>}0$) and shows TCA viability is stable within $0.4$ points across $\kappa\in\{0.00,0.05,0.10\}$. A full graph-structured extension with per-edge latency remains future work.
Synthetic environments. The validation environments use categorical beliefs with Dirichlet-likelihood updates, which admit closed-form Bayesian inference. Agents operating on learned or implicit belief states, where the belief distribution must be approximated from model outputs or sampled posteriors, would require an extension of the belief update mechanism. Section 4.2 outlines one possible bridge, but empirical validation on such systems remains future work.
Single-agent. The current framework models one agent interacting with shared resources. The congestion term $C_{t}$ hints at multi-agent interference, but game-theoretic coupling (multiple agents simultaneously depleting the same tool pool) is not modeled.
5.5 Conclusion
The Triadic Cognitive Architecture offers a unifying perspective: robust autonomy requires explicit coupling of structured space, temporal dynamics, and epistemic truth maintenance. By casting deliberation as a constrained physical trajectory, TCA provides a principled framework for bounded reasoning in tool-using agents. Across two structurally distinct environments, TCA yields consistent 30+30{+}-point improvements in the resource metric over greedy baselines, while preserving accuracy. Together with the purposive stopping baselines and continuation-value analysis, these results support cognitive friction pricing as a domain-independent principle for bounded agent reasoning.
5.6 Broader Impact & Significance
TCA addresses a principled gap in autonomous agent design: the absence of a formal accounting for information-acquisition costs when tool queries are not free, instantaneous, or equally reliable. By formalizing Cognitive Friction and deriving a principled stopping boundary via HJB optimal stopping, approximated through computable rollout-based VOI, the framework transforms deliberation from an unconstrained resource into a measurable, optimizable quantity. The framework is domain-agnostic by construction: the same cost structure applies wherever an agent selects among costly, heterogeneous information sources under time pressure. The two validation environments (emergency medical triage and network security incident response) span distinct urgency regimes and tool profiles, providing controlled evidence that the principle generalizes beyond a single domain. We view TCA as a formal foundation that can be instantiated across a range of agent architectures and deployment settings, with extensions to learned or implicit belief states identified as important future work.
Appendix A Proof sketch for Theorem 1 (idealized continuous-time model)
This appendix provides a proof sketch for the idealized continuous-time formulation in Section 3 (Wonham filtering and HJB-inspired optimal stopping). It does not apply directly to the discrete EMDG controller, which uses rollout-based VOI estimation and a myopic net-utility stopping rule (Section 3.4).
We restate the core claim for the idealized model: under standard regularity conditions, the triadic cognitive control problem admits an optimal stopping time characterized by an HJB free boundary.
Step 1: Regularity of the filtering dynamics and Markov state.
Let the observation process be governed by
$dY_{t}=h(\theta,u_{t})\,dt+\sigma\,dW_{t},$
with $h:\Theta\times\mathcal{A}\to\mathbb{R}^{d}$ bounded and $\sigma>0$ (constant scalar noise). The congestion state evolves as $\dot{C}_{t}=\Omega(u_{t})$ with $\Omega$ bounded. Since $\sigma>0$ is a fixed positive scalar, the noise covariance is $\sigma^{2}I_{d}$, which is uniformly non-degenerate with constant $\varepsilon=\sigma^{2}$. Consequently, given a fixed control $u_{t}$, the innovation process $d\nu_{t}=dY_{t}-\bar{h}(p_{t},u_{t})\,dt$ is a Brownian motion with respect to the observation filtration $\mathcal{F}_{t}^{Y}$, and the associated Wonham (finite-state Kushner–Stratonovich) filter is well-posed; see, e.g., kushner1964differential ; bain2009fundamentals . In particular, the controlled state $(p_{t},C_{t})$ is Markov with respect to $\mathcal{F}_{t}^{Y}$.
Define $\bar{h}(p_{t},u_{t})=\int_{\Theta}h(\theta,u_{t})\,p_{t}(\theta)\,d\theta$ and the innovation
$d\nu_{t}=dY_{t}-\bar{h}(p_{t},u_{t})\,dt.$
A schematic form of the Kushner–Stratonovich update is
[…]
Step 2: Value function and dynamic programming.
The agent seeks to maximize expected information gain minus spatio-temporal cost. For a stopping time $\tau\geq t$, write
$J(\tau)=\mathbb{E}\left[\int_{t}^{\tau}f(p_{s},C_{s},s,u_{s})\,ds\right],$
where $f(p,C,t,u)=\frac{1}{2\sigma^{2}}\mathbb{E}_{p}[\|h(\theta,u)-\bar{h}(p,u)\|^{2}]-\lambda_{S}(C+\Omega(u))-\beta t$ is the instantaneous net reward, consistent with Section 3. Let
$V(p,C,t)=\sup_{\tau\geq t}J(\tau)$
denote the value function. Assuming sufficient regularity of $V$ (or working in the viscosity-solution framework), standard dynamic programming arguments for controlled Markov processes yield a variational inequality characterization shiryaev2007optimal ; oksendal2003stochastic .
Step 3: HJB variational inequality and the free boundary.
By optimal stopping theory, the value function satisfies a variational inequality of the form
$\max\Big\{\partial_{t}V(p,C,t)+\sup_{u\in\mathcal{A}}\big[\mathcal{L}^{u}V(p,C,t)+f(p,C,t,u)\big],\;-V(p,C,t)\Big\}=0,$
in an appropriate classical or viscosity sense shiryaev2007optimal ; oksendal2003stochastic , where $\mathcal{L}^{u}$ is the controlled infinitesimal generator for the Markov state $(p_{t},C_{t})$ as in Section 3. Because $f$ is continuous and bounded under the assumed cost structure, $V$ is continuous under standard assumptions. Define the stopping region as the closed set
$D=\{(p,C,t):V(p,C,t)=0\},$
and the continuation region as the open set $\{V>0\}$. Then the first hitting time of the stopping region,
$T^{*}=\inf\{s\geq t:(p_{s},C_{s},s)\in D\},$
is an optimal stopping time; moreover, it is the minimal optimal stopping time in the usual sense for such problems shiryaev2007optimal . This completes the proof sketch.
Appendix B Action-space scaling: TCA with $|\mathcal{A}|\in\{5,10,20\}$
To assess whether TCA’s advantage over greedy ReAct is specific to the 2-action EMDG setting or generalizes to larger and more varied action spaces, we run a synthetic scaling sweep. For each $|\mathcal{A}|\in\{5,10,20\}$ we generate $N_{\text{cfg}}{=}30$ random tool configurations using a fixed seed, each with independently drawn latency $\tau(a)\in\{3,5,8,10,15,20,30,45\}$ (steps) and information gain $g_{a}\sim\mathcal{U}[0.2,1.5]$ drawn independently, so high-$\tau$ tools can have high information gain and vice versa, forcing the optimizer to evaluate genuine gain-versus-cost tradeoffs. Each configuration is evaluated over $N_{\text{seed}}{=}10$ independent trajectory seeds. All other parameters ($\alpha{=}0.01$, $\beta{=}0.5$, $\lambda_{S}{=}0.8$) are held at their EMDG defaults.
Non-trivial step-0 selection.
For each run we record whether TCA’s first action is the minimum-latency tool in the configuration. Selecting the cheapest tool regardless of gain would be trivial cost-minimization; selecting a higher-$\tau$ tool when its information gain justifies the cost demonstrates the genuine cost-aware optimization claimed by TCA.
Results.
Table 7 reports terminal statistics. The fraction of (configuration, seed) pairs in which TCA selects a non-minimum-latency tool at step 0 rises monotonically: $76\%$ at $|\mathcal{A}|{=}5$, $85\%$ at $|\mathcal{A}|{=}10$, and $90\%$ at $|\mathcal{A}|{=}20$. (For comparison, this fraction is $0\%$ in the 2-action EMDG, where Hematology Lab is always preferred over MRI; see Table 1.) TCA terminal viability remains stable at 91.891.893.993.9 across all sizes, fully consistent with the main-text EMDG result of $93.03\pm 0.19$. TCA consistently outperforms ReAct (gap $\geq+9.7$ viability points; accuracy $1.00$ throughout).
The absolute gap is smaller than in the 2-action EMDG ($\approx$10–15 vs. 36 points), which is expected: in random tool configurations, the greedy max-gain rule occasionally selects a low-latency tool when the highest-gain tool happens to have moderate latency, reducing, but never eliminating, TCA’s advantage. Crucially, TCA dominates in every configuration tested and its viability is stable across $|\mathcal{A}|$, confirming that the controller scales gracefully and performs genuine cost-aware optimization rather than a degenerate “always pick cheapest” rule. We view this sweep as a robustness check that the policy does not collapse to trivial cost minimization; it is not intended as an exhaustive scaling benchmark.
Table 7: Action-space scaling ($N_{\text{cfg}}{=}30$ random tool configurations per $|\mathcal{A}|$, $N_{\text{seed}}{=}10$ seeds per configuration). Non-trivial step-0: fraction of (config, seed) pairs where TCA selects a tool with $\tau>\min_{a}\tau(a)$. Viability: mean $\pm$ 95% CI. For $|\mathcal{A}|{=}2$, see Table 1.
$|\mathcal{A}|$ Non-trivial step-0 TCA Viability ReAct Viability Gap Accuracy
55 $76\%$ $91.78\pm 0.50$ $81.94\pm 1.52$ +9.8+9.8 $1.00$
$10$ $85\%$ $92.62\pm 0.49$ $82.89\pm 1.35$ +9.7+9.7 $1.00$
2020 $90\%$ $93.85\pm 0.38$ $78.84\pm 1.61$ +15.0+15.0 $1.00$
Appendix C Discrete Controller Pseudocode
For completeness, we summarize the discrete controller used in EMDG and NSTG. The box below is intentionally faithful to Section 3.4: it is the rollout-VOI, myopic net-utility controller actually analyzed in the main text, not a stronger planner.
Table 8: Algorithm 1: Discrete TCA controller (rollout VOI + myopic stopping).
Inputs: current belief $b_{t}$, congestion $C_{t}$, time $t$, action set $\mathcal{A}$, rollout count $K$, cost parameters $\alpha,\beta,\lambda_{S}$.
For each action $a\in\mathcal{A}$:
1. Clone the current belief state $K$ times. 2. For each clone, sample a tool observation under action $a$, perform the Bayesian update, and record the entropy drop $H(b_{t})-H(b_{t+1}^{(k)}(a))$. 3. Estimate rollout value-of-information: $\widehat{\Delta H}(a\mid b_{t})=\frac{1}{K}\sum_{k=1}^{K}\Big(H(b_{t})-H(b_{t+1}^{(k)}(a))\Big).$ 4. Compute net utility: $U(a;b_{t},t,C_{t})=\widehat{\Delta H}(a\mid b_{t})-\alpha\Big(\lambda_{S}(C_{t}+\Omega(a))+\beta\,(t+\tau(a))\Big).$ Select $a^{*}=\arg\max_{a\in\mathcal{A}}U(a;b_{t},t,C_{t})$.
If $U(a^{*};b_{t},t,C_{t})\leq 0$, return Stop.
Else execute $a^{*}$, observe the realized outcome, update the belief by Bayes’ rule, and advance the environment state:
$C_{t+1}=C_{t}+\Omega(a^{*}),\qquad t_{+}=t+\tau(a^{*}).$ Repeat until the stop condition is met.